OpenStudy (anonymous):

e^2x - 2e^x = +8 then x = ???

OpenStudy (baldymcgee6):

Think of this as a quadratic

OpenStudy (baldymcgee6):

Just move the eight to the other side, (e^x)^2 -2e^x - 8 = 0

OpenStudy (baldymcgee6):

or how about let z = e^x (z)^2 -2z - 8 = 0

OpenStudy (raden):

or factor direct can be (e^x - 4)(e^x + 2) = 0

OpenStudy (raden):

now, for zeros all terms take e^x-4=0 or e^x+2=0, solve for x's!